OpenStudy (gmail.comlover):

if f(x) = 0.5 (1-x), what is the value of f(-1)? PLEASE HELP

OpenStudy (michele_laino):

it is simple, we have to set \(x=-1\), so we can write this: \[f\left( 1 \right){\text{ }} = {\text{ }}0.5{\text{ }}\left( {1 - \left( { - 1} \right)} \right) = ...?\]

OpenStudy (michele_laino):

oops... \[f\left( { - 1} \right){\text{ }} = {\text{ }}0.5{\text{ }}\left( {1 - \left( { - 1} \right)} \right) = ...?\]

OpenStudy (michele_laino):

:)